OpenStudy (ny,ny):

how would i get a by itself in -1/4a=2/5

OpenStudy (anonymous):

divided both sides by -1/4

OpenStudy (anonymous):

which is the same as multiplying -4 on both side

OpenStudy (ny,ny):

so multiply -4 with 2/5?

OpenStudy (anonymous):

yes

OpenStudy (ny,ny):

-1.6. is this what a is equal to?

OpenStudy (anonymous):

which leads to a = -8/5

OpenStudy (anonymous):

yes

OpenStudy (ny,ny):

thanks

OpenStudy (ny,ny):

do i follow the same steps for the problem: 12=9/5y

OpenStudy (anonymous):

multiply both sides by 5/9

OpenStudy (ny,ny):

u flipped it. is that necessary?

OpenStudy (anonymous):

yes so you get y = 20/3